OpenStudy (ram231):

How do you find diagonals of a rectangle? are they the lines that go from corner to corner across the middle?

OpenStudy (landon34):

Yes I think that is true.

OpenStudy (anonymous):

You can use pythagorean theorem And yes that is true

OpenStudy (anonymous):



OpenStudy (ram231):

is the pythagorean theorem basically like a2+b2=c2? And i have the length of 36 and diagonals of 39.

OpenStudy (anonymous):

If you know the 2 legs you know the diagonals.

OpenStudy (anonymous):

So to do that problem you would rearrange the equation to c^2-a^2=b^2

OpenStudy (ram231):

how do I find the area though? thats the confusing part

OpenStudy (anonymous):

Once you get the other leg it's just length x width

OpenStudy (ram231):

I'm confused. so I have the length of 36 and the hypotenuse of 39. so I just do 39^2 - 36^2 and that answer would be the other length and then I would just have to do length x width?

OpenStudy (anonymous):

Well after you do 39^2-36^2 you would take the square root of that and that would give you the leg and the rest is correct

OpenStudy (ram231):

ok. ill try it

OpenStudy (ram231):

I still don't get it....

OpenStudy (ram231):

@zepdrix

OpenStudy (anonymous):

because \[39^2-36^2=c^2\rightarrow225=c^2\] so to get c you would have to take the square root

OpenStudy (ram231):

okay is it 540?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

Sorry that equation was a bit confusing should've said b instead of c

OpenStudy (anonymous):

Make sure you put 540 units squared

OpenStudy (ram231):

well it only gives me a choice of 540. I can't square it.